OpenStudy (anonymous):

Can somepne please help me Solve for x: −3|x + 7| = −12

OpenStudy (iwillrektyou):

I will try to help, but it may take me a bit of time to remember. This wasn't my favorite topic in algebra, but I will try my best.

OpenStudy (iwillrektyou):

What are the choices (if any)?

OpenStudy (anonymous):

x = −13 and x = −3 x = 3 and x = −3 x = 3 and x = 13 No solution

OpenStudy (iwillrektyou):

Every absolute equation has 2 solutions.

pooja195 (pooja195):

Divide both sides by 3 −3|x + 7| /3= −12/3 |x+7|=4 We know either x+7=4or x+7=−4

OpenStudy (anonymous):

Okay But I think it is no solution

pooja195 (pooja195):

x+7=4(Possibility 1) x+7=−4(Possibility 2) Solve both for x this one does have 2 solutions

OpenStudy (iwillrektyou):

For possibility 1, x + 7 = 4 -7 - 7 x = -3

pooja195 (pooja195):

Can you let the asker solve it please?

OpenStudy (anonymous):

So what do you think that it's the second choice?

OpenStudy (iwillrektyou):

Solve possibility 2 for x: x+7=−4

OpenStudy (anonymous):

So it's the first one becuase x+7+-4 is -13

OpenStudy (iwillrektyou):

There are 2 solutions for each absolute value equation. Unless one doesn't have a solution, both are correct.

OpenStudy (anonymous):

All the choices have 2 answers in them so I mean is it the first choice

OpenStudy (iwillrektyou):

Oh, sorry for misunderstanding. then yes, it is.

OpenStudy (anonymous):

thank you so so much

OpenStudy (iwillrektyou):

Np! Some of the credit also goes to pooja195

OpenStudy (anonymous):

I know thanks to all of y'll

pooja195 (pooja195):

wow..